# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v3

@@ Appendix A Feature Extraction
Figure 9: The hand-crafted convolutional kernels of the feature extractor.
The stage $S1$ serves to extract initial binary features from the images. The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 12).
@@ Appendix B Normalization
The membrane potential in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{M}(S2)}$ to the interval $(0,1)$.
@@
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (21)
According to this formula, $\boldsymbol{a}^{\mathcal{M}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{M}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i^{\prime},t}},0\right)$ (22)
Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$. This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values. This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
@@ Appendix C Single Weight-Kernel Implementation
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (23)
@@ Appendix D Dataset
The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles). The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$. Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location. To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures. These images include line drawings of digits 00-$9$, characters, and objects such as a house, a star, a wheel, a clock, etc. Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
Appendix E Parameters
We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs). The model uses $\kappa=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix F Autoencoder Parameter
@@
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels. The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
To provide a robust comparison, we increase the number of training samples for the autoencoder. Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch. This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction. The autoencoder is trained with a batch size of $512$.
@@ Appendix G Quantifying Noise Filtration
In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$. Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 10).

### v4

@@ Appendix A Feature Extraction
Figure 9: Features: The hand-crafted convolutional kernels of the feature extractor.
The stage $S1$ serves to extract initial binary features from the images. The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $\times 5$ 5 as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of 111 (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size (1×32×32)(1$\times$ 32$\times$ 32)( 1 × 32 × 32 ) and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size (4×32×32)(4$\times$ 32$\times$ 32)( 4 × 32 × 32 ). The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 18).
@@ Appendix B Normalization
The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval (0,1)(0,1)( 0 , 1 ).
@@
$\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{I}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (31)
According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2- 1 / 2. (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{I}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{I}(S2)}_{c,i^{\prime},t}},0\right)$ (32)
Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$. This normalization confines activations to the interval [0,1][0,1][ 0 , 1 ], where neurons receiving maximal support are mapped to 111, while those with less support are mapped to correspondingly smaller values. This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
@@ Appendix C Single Weight-Kernel Implementation
$\boldsymbol{a}^{\mathcal{I}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (33)
@@ Appendix D Dataset
The dataset used for training comprises binary images measuring ($\times 32$ 32) pixels, each sample depicting a straight line going through the image center, starting and ending 222 pixels from the image boundary (leading to 595959 distinct lines of different angles). The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate 300300300 image instances (each line about 555 times). During evaluation, we sample each of the 595959 distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
[…]
To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($\times 32$ 32) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
Additionally, we generate binary images measuring ($\times 64$ 64) pixels by hand that represent more complex structures. These images include line drawings of digits 0-999, characters, and objects such as a house, a star, a wheel, a clock, etc. Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
Appendix E Training Parameters
We train the model with a learning rate of $\eta=0.2$ for 100100100 training cycles (epochs). The model uses $\kappa=10$ competitive neurons with recurrent connections spanning a kernel of $\times 11$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of competitive channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix F Autoencoder Training
@@
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of 333 and a stride of 222, with progressively increasing channel counts of 323232, 646464, 128128128, and 256256256 channels. The decoder mirrors this structure, using transposed convolutional layers with 256256256, 128128128, 646464, and 323232 channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with 777777777k parameters, utilizing up to 256256256 channels, whereas the CNA model operates with only 444444 channels.
The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ), and a weight decay of $1\times 10^{-8}$.
To provide a robust comparison, we increase the number of training samples for the autoencoder. Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for 200200200 epochs, using 10,00010,00010 , 000 images per epoch. This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction. The autoencoder is trained with a batch size of 512512512.
@@ Appendix G Quantifying Noise Filtration
[…]